OpenStudy (anonymous):

y = 3x - 87 3x + 2y = 21 solve using the substitution method ....?

OpenStudy (anonymous):

Put y = 3x -87 in second eq. 3x +2(3x -87) = 21 find x then use that value of x to find y

OpenStudy (anonymous):

Im confused.. could you explain it a little better?? like in detail..??

OpenStudy (anonymous):

in substitution method, substitute one unknown (x) in place of second unknown (y) you have two equations y = 3x -87 3x +2y = 21 now use value of y from first eq. in second eq 3x + 2(3x-87) = 21 3x +2*3x - 2*87 = 21 3x+6x - 174 = 21 9x = 174 +21 9x = 195 x =195/9 = 21.6 Now put this value of x in first eq. to find y

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

that's not how I did it though..

OpenStudy (anonymous):

getting correct answer is important